OpenStudy (unklerhaukus):

Solve the differential Equation \[\frac{dy}{dx}+\frac{2x−y−4}{ 2x − y + 5}=0\]

OpenStudy (anonymous):

Try writing v = 2x - y. That may simplify by a bunch.

OpenStudy (anonymous):

Then: dv/dx = 2 - dy/dx. Rewrite dy/dx as 2 - dv/dx :-)

OpenStudy (anonymous):

That becomes a separable DE.

OpenStudy (unklerhaukus):

\[2-\frac{\text d v}{\text d x}+\frac{v-4}{v+5}=0\] this looks seperable

OpenStudy (anonymous):

yup :-)

OpenStudy (unklerhaukus):

thankyou bmp

OpenStudy (anonymous):

No problem :-)

OpenStudy (unklerhaukus):

\[\frac{\text d v}{\text d x}=2+\frac {v-4}{v+5}\]\[\frac{\text d v}{\text d x}=\frac{2(v+5)+(v-4)}{v+5}\]\[\frac{\text d v}{\text d x}=\frac {3v+6}{v+5}\]\[\frac{\text d v}{\text d x}=\frac {3(v+2)}{v+5}\]\[\frac{\text d x}{\text d v}=\frac {v+5}{3(v+2)}\]\[\text d x=\frac {v+5}{3(v+2)}\text d v\]\[\int\text d x=\frac 1 3\int\frac {v+5}{v+2}\text d v\] right so far?

OpenStudy (unklerhaukus):

\[x=\frac13 \int\frac {v}{v+2}+\frac {5}{v+2}\text d v\]

OpenStudy (experimentx):

let x = X+h and y = Y+k <--- reduce it into homogeneous differential equation.

OpenStudy (unklerhaukus):

@experimentX are you suggesting that the method i have working for, is not the bet way to answer this question?

OpenStudy (experimentx):

I haven't looked ... i remembered that i solved these kinds of differential equation by reducing it to homogeneous differential equation. 2h - k - 4 =0 2h - k + 5 =0 Oo ... looks like your method is right!!!

OpenStudy (unklerhaukus):

it could be that both methods work but i might try the first method to the end first

OpenStudy (unklerhaukus):

\[x=\frac13 \int\frac {v}{v+2}+\frac {5}{v+2}\text d v\]\[x=\frac{1}{3} \left( v-2\log(v+2)+5\log(v+2)\right) + c_1\]

OpenStudy (unklerhaukus):

\[x=\frac{1}{3} \left( v+3\log(v+2)\right) + c_1\]\[x= \frac{(2x-y)}{3}+\log((2x-y)+2) + c_1\]\[x= \frac{(2x-y)}{3}+\log(2x-y+2) + c_1\]

OpenStudy (unklerhaukus):

now what should i do?

OpenStudy (anonymous):

\[x= \frac{(2x-y)}{3}+\log(2x-y+2) + c_1\] this is the solution.. you cannot do anything beyond this..

OpenStudy (unklerhaukus):

ok cool